OpenStudy (anonymous):

WRITEN RESPONCE----- Bryan has to carry 24 books to a new shelf. If he carries one book at a time, it will take him 24 trips. If he carries 3 books at a time, how many trips will it take? Does this relationship represent direct or inverse variation? Using the appropriate formula, find k. Show your work and answer the original question.

OpenStudy (anonymous):

8 trips...?

OpenStudy (anonymous):

This is inverse relation, y=k/x Books trips 1 - 24 3 - x 3x=24X1 x= 24/3 x=8